If f(x)= x^2 + 7, what is the equation for f^-1(x)?

Answer by stanbon(75887)

If f(x)= x^2 + 7, what is the equation for f^-1(x)?
----
Notice that x is 1st of all squared, then 7 is added.
----
The inverse requires that you reverse the process:
1st you subtract 7 ; then you take the square root.
----
Ans: f^-1(x) = sqrt(x-7)
==============================
Process:
1st interchange x and y to get:
x = y^2 + 7
2nd solve for "y":
y^2 = x-7
f^-1(x) = sqrt(x-7)
===================
cheers,
Stan H.